combining 0.385 mol of feo3 with excess carbon produced 16.5g of fe.
what is the actual yield of iron in moles?
what was the theoretical yield of iron in moles?
what was the percent yield?
1
Expert's answer
2018-02-14T06:15:51-0500
.n (FeO3)=0.385 mol
.m (Fe)=16.5 g
.n(Fe)theor=?
%=?
FeO3+ 3C=Fe+3CO
.n (Fe)theor= n (FeO3)=0.385 mol
Amount of substance of iron n (Fe)=m/Ar= 16.5/56=0.295 mol
%= n(Fe)/ n (Fe) theor= 0.295/0.385 *100 %= 76.62 %